Please solve using Kirchhoffs rules to determine the voltage and current through each resistor.

4ΚΩ V2 6ον 80ν ΟΙ R2 >3kΩ ΜΜΜ 2kΩ.
Answer #1

u 4k likti I sou zou 2 km Apply KCL at We => Z2 = 2, +32 by ohms law I, = 70-us. 1= som en 13 = ve 34 ; VC-602- : 70-u 2k so

If you have any doubt please make a comment. Thank you